The position of a particle is given by r(t)= t2i + e2tj + In(t+1)k. Find the minimum velocity of the particle


1
Expert's answer
2021-08-20T09:00:56-0400
"r(t)=t^2i + e^{2t}j + \ln(t+1)k"

"v(t)=r'(t)=2ti +2 e^{2t}j + \dfrac{1}{1+t} k"

Let "f(t)=(|v(t)|)^2=(2t)^2+(2e^{2t})^2+(\dfrac{1}{1+t})^2"


"f'(t)=8t+16e^{4t}-\dfrac{2}{(1+t)^3}\geq0+16-2>0, t\geq0"

Therefore the function "f(t)" increases for "t>0."

Since "t\geq0" (time), then the minimum value of "f" for "t\geq 0" is


"f(0)=(2(0))^2+(2e^{2(0)})^2+(\dfrac{1}{1+0})^2=5"

The minimum velocity of the particle is "\sqrt{5}."